OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 8x + y = 10 6x + 5y = –1

OpenStudy (firejay5):

Your first step is to solve the first equation 8x + y = 10, -8x to both sides and your new equation should be y = 10 - 8x

OpenStudy (firejay5):

substitute 10 - 8x into the second equation for y in order to solve for x. 6x + 5(10 - 8x) = -1. Then use Distributive Property and the equation will look like this: ---> 6x + 50 - 40x = -1

OpenStudy (firejay5):

-50 from both sides and subtract 6x - 40x = -34x and -1 - (-50) = -51 your problem should be -34x = -51; divide by -34 and you get x = 1.5 or 3/2. Plug 3/2 in either equation to solve for y.

OpenStudy (firejay5):

8(3/2) + y = 10 12+ y = 10 -12 -12 y = -2

OpenStudy (firejay5):

I hope this helped you out @slkennedy

OpenStudy (anonymous):

Yes thank you, you broke it down in easy to understand terms. thanks